If X has a Poisson distribution with parameter λ, and if
P(X=0)=0.2.
Evaluate
P(X>2).
1
Expert's answer
2020-10-05T17:41:47-0400

Poisson distribution has following PDF:

"P(X=k) = \frac{\lambda^k}{k!} e^{-\lambda}"

We can use given value "P(X=0)=0.2" to find the value of "\lambda".

"P(X=k) = \frac{1}{1} e^{-\lambda} = 0.2"

"\lambda = - \ln{0.2=1.609}", and, since the coefficient "e^{-\lambda}" is the same for every X=k, we can use directly "e^{-\lambda}=0.2".


"P(X>2) = 1 - P(X \leq 2) = 1 - P(X=0) - P(X=1) - P(X=2)"

"P(X=0)=0.2"

"P(X=1) = \frac{1.609}{1} 0.2 =0.322"

"P(X=2) = \frac{1.609^2}{2 \cdot 1} 0.2=0.259"

So, "P(X>2) = 1 -0.2-0.322-0.259 = 0.219"

Answer: "P(X>2)=0.219"